• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name moffit

    Log entry time 08:02:12 on October 5,2002

    Entry number 87351

    keyword=shift summary

    00:00: MCC still working to restore beam.
    00:58: MCC called and let us know that their oncall Rad expert is under the impression that there could be a target misalignment causing the ion chamber trips. JP paged. JP suggested an access to take a visual inspection of the target ladder. Asked MCC for an access.
    01:20: Hall is now in controlled access. Ron is taking a look.
    01:32: Target moved to optics position.
    01:42: Mike Seely Paged.
    02:03: Target moved to Thin Carbon position
    02:12: Hall returns beam permit.
    02:17: Left Q1 appears to have tripped. Asked MCC for another access to reset both left and right Q1s.
    02:34: Ron has gone down to reset the Q1s. Cryo Coordinator is onsite and is investigating ESR fluctuations. Ed Folts suggests having the cryo coord. raise the lower limit of the fill valve by 5%. Cryo Coord. will do that for left Q1 by 10%.
    02:53: Ron has reset left and right Q1s.
    03:03: Asked MCC for beam on Thin Carbon target.
    03:17: MCC sends pulsed beam to look at orbits.
    03:52: MCC reports lots of BLM trips. Still working on it.
    04:01: MCC done with steering... will send 10uA of CW.
    04:04: MCC cannot deliver beam. Have paged an expert.
    05:22: MCC sends 5uA of pulsed beam to the Hall-A dump
    06:20: MCC awaiting optics expert to work on orbits (on-site instead of on-phone)
    07:05: MCC says they've made some changes at the injector and are going to send CW to the Hall-A dump
    07:28: Target moved to empty position so MCC can check beam stability
    07:56: MCC says they can send a steady 20uA. Move target to Thick Carbon position.

    BTW: Hall Access/Entry Sheets appear to be missing from the counting house. Were they left in the Hall?